OpenStudy (anonymous):

36a^2 - 9 factor the polynomial

OpenStudy (mathstudent55):

Have you heard of the difference of two squares?

jimthompson5910 (jim_thompson5910):

Use the idea that x^2 - y^2 = (x-y)(x+y)

OpenStudy (mathstudent55):

The difference of two squares factors as follows: x^2 - y^2 = (x + y)(x - y) 36a^2 is the square of 6a 9 is the square of 3 In your case, x = 6a, and y = 3